### v1

@@ Abstract
Spiking Neural Networks (SNNs) are highly regarded for their energy efficiency, inherent activation sparsity, and suitability for real-time processing in edge devices. However, most current SNN methods adopt architectures resembling traditional artificial neural networks (ANNs), leading to suboptimal performance when applied to SNNs. While SNNs excel in energy efficiency, they have been associated with lower accuracy levels than traditional ANNs when utilizing conventional architectures. In response, in this work we present LightSNN, a rapid and efficient Neural Network Architecture Search (NAS) technique specifically tailored for SNNs that autonomously leverages the most suitable architecture, striking a good balance between accuracy and efficiency by enforcing sparsity. Based on the spiking NAS network (SNASNet) framework, a cell-based search space including backward connections is utilized to build our training-free pruning-based NAS mechanism. Our technique assesses diverse spike activation patterns across different data samples using a sparsity-aware Hamming distance fitness evaluation. Thorough experiments are conducted on both static (CIFAR10 and CIFAR100) and neuromorphic datasets (DVS128-Gesture). Our LightSNN model achieves state-of-the-art results on CIFAR10 and CIFAR100, improves performance on DVS128Gesture by 4.49%, and significantly reduces search time most notably offering a 98×98$\times$98 × speedup over SNASNet and running 30% faster than the best existing method on DVS128Gesture.
@@ IV Experiments and evaluation
Table 2: Performance comparison of NAS methods for the CIFAR10 dataset.
@@ IV-B Hyperparameters
For the search phase, weights are randomly initialized with the Kaiming Initialization [39], and the search batch size is set to $32$. Different search batch sizes could have been investigated but were not tested in this work. For a $300$-epoch training phase, the surrogate gradient method is run to enable backpropagation in SNNs [14], the batch size was set to $64$, with $0.2$ as the learning rate with a cosine-annealing learning rate schedule. We used the vanilla SGD optimizer with a momentum of $0.9$ and weight decay of $0.0005$. The algorithm was implemented using Pytorch and SpikingJelly libraries [^1] Code will be available upon completion of the review process. and executed on an Nvidia A40 GPU.
@@ IV-C Results
To assess the effectiveness of our work, we consider three criteria: accuracy, search time, and spiking activity rate (SAR). The accuracy informs us about how well the model could learn and generalize to test data. The search time evaluates the efficiency of the proposed NAS framework and confirms its rapidity. Finally, the SAR quantifies the frequency of spike generation within the network. Specifically, the SAR is used as a proxy metric for sparsity – which generally results in lightweight and energy efficient SNN models [37]. The SAR metric is here evaluated by dividing the total number of spikes by the total number of neurons multiplied by the total number of time steps.
LightSNN employs a reduced SNASNet search space and the training-free, pruning-by-importance algorithm.
For the accuracy, as shown in Tab. 2, LighSNN performs competitively, reaching $93\%$ on CIFAR10, slightly below SpikeNas at $93.18\%$, but higher than SNASNet at $91.83\%$. On CIFAR100, LightSNN reaches $70.44\%$, much higher than SpikeNas, which scored $45.77\%$, and competing well with SNASNet, which reached $72.36\%$. On DVS128Gesture, LightSNN reaches the maximum accuracy of $94.44\%$ surpassing both SNASNet ($89.93\%$) and SpikeNas ($87.84\%$).
In terms of search time, LightSNN greatly outperforms SNASNet, but it is slower compared to SpikeNas on both CIFAR10 and CIFAR100. On CIFAR10, the search of LightSNN completes in $2$m $44$s, much faster compared with $2$h $49$m of SNASNet, but slower compared with SpikeNas which uses $29$s. On CIFAR100, our LightSNN completes the search in $5$ minutes $58$ seconds, still faster than SNASNet’s $2$ hours $2$ minutes, yet slower compared to SpikeNas which takes $3$ minutes $3$ seconds. However, on the DVS128Gesture, LightSNN finishes the search in $6$ minutes $54$ seconds, beating the previous two at $11$ hours $29$ minutes and $9$ minutes $55$ seconds, respectively, for SNASNet and SpikeNas, demonstrating scalability to more complex and dynamic datasets.
LightSNN reaches competitive results concerning SAR across all datasets. For the DVS128Gesture, the SAR is relatively low at $0.07$, outperforming SNASNet (at $0.13$) and coming close to SpikeNas (at $0.05$). On both CIFAR10 and CIFAR100, LightSNN performs close to SpikeNas scoring a NAS of $0.09$ and $0.13$, for CIFAR10 and CIFAR100, respectively. This shows that LightSNN performance is overall very satisfactory, showing the best results for dynamic datasets and competitive results in all other cases, for all the considered metrics.

### v2

@@ Abstract
Spiking Neural Networks (SNNs) are highly regarded for their energy efficiency, inherent activation sparsity, and suitability for real-time processing in edge devices. However, most current SNN methods adopt architectures resembling traditional artificial neural networks (ANNs), leading to suboptimal performance when applied to SNNs. While SNNs excel in energy efficiency, they have been associated with lower accuracy levels than traditional ANNs when utilizing conventional architectures. In response, in this work we present LightSNN, a rapid and efficient Neural Network Architecture Search (NAS) technique specifically tailored for SNNs that autonomously leverages the most suitable architecture, striking a good balance between accuracy and efficiency by enforcing sparsity. Based on the spiking NAS network (SNASNet) framework, a cell-based search space including backward connections is utilized to build our training-free pruning-based NAS mechanism. Our technique assesses diverse spike activation patterns across different data samples using a sparsity-aware Hamming distance fitness evaluation. Thorough experiments are conducted on both static (CIFAR10 and CIFAR100) and neuromorphic datasets (DVS128-Gesture). Our LightSNN model achieves state-of-the-art results on CIFAR10 and CIFAR100, improves performance on DVS128Gesture by 4.49%, and significantly reduces search time most notably offering a 98×98$\times$98 × speedup over SNASNet and running 30% faster than the best existing method on DVS128Gesture. Code is available on Github at: https://github.com/YesmineAbdennadher/LightSNN.
@@ IV Experiments and evaluation
Table 2: Performance comparison of NAS methods for different datasets.
@@ IV-B Hyperparameters
For the search phase, weights are randomly initialized with the Kaiming Initialization [39], and the search batch size is set to $32$. Different search batch sizes could have been investigated but were not tested in this work. For a $300$-epoch training phase, the surrogate gradient method is run to enable backpropagation in SNNs [14], the batch size was set to $64$, with $0.2$ as the learning rate with a cosine-annealing learning rate schedule. We used the vanilla SGD optimizer with a momentum of $0.9$ and weight decay of $0.0005$. The algorithm was implemented using Pytorch and SpikingJelly libraries and executed on an Nvidia A40 GPU.
@@ IV-C Results
To assess the effectiveness of our work, we consider three criteria: accuracy, search time, and spiking activity rate (SAR). The accuracy informs us about how well the model could learn and generalize to test data. The search time evaluates the efficiency of the proposed NAS framework and confirms its rapidity. Finally, the SAR quantifies the frequency of spike generation within the network, directly corresponding to the energy consumption of an SNN  [40]. Specifically, the SAR is used as a proxy metric for sparsity, which generally results in lightweight and energy-efficient SNN models [37]. The SAR metric is here evaluated by dividing the total number of spikes by the total number of neurons multiplied by the total number of time steps.
Compared to SNASNet, lightSNN achieves significantly higher accuracy on CIFAR-10 and DVS128Gesture and comparable accuracy on CIFAR-100, while substantially reducing search time across all datasets. It also yields lower sparsity on CIFAR-10 and DVS128Gesture and maintains similar sparsity on CIFAR-100. Although lightSNN incurs a modest runtime penalty relative to SpikeNAS on CIFAR-10 and CIFAR100, due to its search method, it achieves substantially higher accuracy on CIFAR-100 and nearly matches SpikeNAS on CIFAR10 at equivalent sparsity. Conversely, on DVS128Gesture, lightSNN not only outperforms SpikeNAS but also runs faster while maintaining similar sparsity.
We highlight our model’s exceptional performance on the DVS128Gesture benchmark, underscoring its suitability as a prime candidate for event-based data classification.